Fermi National Laboratory


Accelerator Update for June 20, 2005 - June 22, 2005

Monday June 20
The day shift began with Operations monitoring store 4219, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE, NuMI, and Meson MCenter experiments taking beam.

At 9:30 AM, Operations reported a small LCW leak from under a Main Injector (MI) power supply (MI-62). An expert investigated and stated that the power supply could wait for repairs until the next shutdown.

At 10:54 AM, a LCW expert from CUB reported that they lost a LCW pump for the Booster 95 system. He switched to pump number 2. At 10:59 AM, the Tevatron Electron Lens (TEL) tripped off. The expert said it had lost its cooling. At 11:15 AM, Operations reported that the Booster RF had all tripped off due to no cooling. Operators put the RF in standby.

Operations began at 11:53 AM preparing MI for an access to investigate an interlock problem. Interlock technicians found and fixed a problem in a reset box.

A Booster expert began a study period at 12:19 PM.

MI came out of access around 3:10 PM. Operations reestablished the MI ramp at 3:54 PM.

Operations established beam to MiniBooNE and NuMI at 4:07 PM.

Pbar resumed stacking at 4:19 PM.

Operations established beam to the Meson experiment MTest at 4:42 PM.

At 5:04 PM, the NuMI beam permit tripped off due to what an expert reported as a false indication from a LCW skid. Experts continued their investigations.

Operations put the Recycler into shot setup at 7:14 PM to stash antiprotons.

The Recycler stash ended and Operations resumed sending beam to MCenter and MTest at 7:50 PM.

TeV experts began an end-of-store study at 9:07 PM.

Operations terminated store 4119 at 9:57 PM.

TeV experts began a chromaticity study at 10:23 PM.

Tuesday June 21
The midnight shift began with Operations monitoring the Recycler stash, with TeV conducting octupole studies, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

MTest experimenters said they were done with the beam at 1:08 AM.

TeV studies ended at 1:56 AM. Operators prepared the Collider Detector at Fermilab (CDF) for a collision hall access.

CDF came out of access at 3:26 AM.

Operations put the TeV through a dry squeeze at 3:50 AM.

Operations put the TeV into shot setup at 4:14 AM.

Operations established combination store 4221 at 6:44 AM with an initial luminosity of 88.94E30. Pbar resumed stacking.

MCenter resumed taking beam at 6:52 AM.

At 7:23 AM, a MI RF station (MIRF5) tripped off for the first of ten times in about ninety minutes. It reset okay each time.

Operations prepared Recycler to stash antiprotons 11:33 AM. The stash was completed by 1 PM. Pbar resumed stacking.

MCenter resumed taking beam at 1:22 PM and MTest took beam at 1:39 PM.

Pbar experts began a stacking study at 1:59 PM.

At 4:22 PM, Operations reported that NuMI had tripped off for the first of a number of times because of beam in its abort gap. Operators traced the problem to the timeline. An expert noticed that one event was too close to another, which was causing power supply problems. Unfortunately, this didnít solve the problem, for NuMI continued to trip. A NuMI expert thought the problem might be related to MI slipstacking.

Pbar experts completed their study at 5:43 PM.

Operations prepared the Recycler to stash antiprotons at 8:09 PM. Operations completed the stash and Pbar resumed stacking at 8:47 PM. MCenter and MTest resumed taking beam.

MTest experimenters reported at 11:19 PM that they were through with the beam for the night.

Wednesday June 22
The midnight shift began with Operations monitoring store 4221 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 5:03 AM, Operations reported a possible power glitch. The Boosterís West Anode Power Supply (WAPS) and the TEL tripped off. Also, some Accumulator cooling was lost. Everything reset within a few minutes. The Duty Electrician investigated at the master substation, but reported at 5:19 AM that he couldnít find a problem.

The Plan for Monday
The plans for today are to allow an end-of-store study and then terminate the store this morning. D-Zero will be allowed an access. The next store will go in early this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/22/2005   email Fermilab